1. What is the angular momentum of a 0.240-kg ball rotating on the end of a thin string in a circle of radius 1.35 m at an angular speed of 15.0 rad/s ?

2. A diver can reduce her moment of inertia by a factor of about 4.0 when changing from the straight position to the tuck position. If she makes 2.0 rotations in 1.5 s when in the tuck position, what is her angular speed (rev/s) when in the straight position?

3. A person of mass 72 kg stands at the center of a rotating merry-go-round platform of radius 3.3 mand moment of inertia 950 kg⋅m2 . The platform rotates without friction with angular velocity 0.85 rad/s . The person walks radially to the edge of the platform. a. Calculate the angular velocity when the person reaches the edge. b. Calculate the rotational kinetic energy of the system of platform plus person before the person's walk.

4. Four identical particles of mass m are mounted at equal intervals on a thin rod of length ℓ and mass M, with one mass at each end of the rod. (a. If the system is rotated with angular velocity ω about an axis perpendicular to the rod through one of the end masses, determine the kinetic energy of the system. (b. If the system is rotated with angular velocity ω about an axis perpendicular to the rod through one of the end masses, determine the angular momentum of the system.
Answer #1

1)

Angular momentum is L = I*w

I is the moment of inertia = m*r^2 = 0.24*1.35^2 = 0.4374 kg-m^2

w = 15 rad/s

then

L = I*w = 0.4374*15 = 6.56 kg-m^2/s